# Disentangling the galactic and intergalactic components in 313 observed Lyman-alpha line profiles between redshift 0 and 5

### v1

@@ 4 Results on observed Lyman-$\alpha$ line profiles
In this section, we apply zELDA’s ANN models to observed ${\rm Ly}\alpha$ line profiles. In Sect. 4.1 we display the predicted ${\rm Ly}\alpha$ line profiles and in Sect. 4.2 we compare the shell model parameters of the different ANN models. Later, in Sect. 4.3 we study zELDA’s accuracy to provide the source redshift in Sect. 4.5. We present our observed $\langle f_{\rm esc}^{4\AA}\rangle$ reconstruction in Sect. 4.6 and we compare it to the global ${\rm Ly}\alpha$ escape fraction in Sect. 4.7 and the intrinsic stacked ${\rm Ly}\alpha$ line profile in Sect. 4.8. We explore the fact that the IGM+z and IGM-z models predict that the ${\rm Ly}\alpha$ line profiles are IGM attenuated at $z\sim 0$ in Sect. 4.9. Finally, we discuss possible biases that might arise from fitting IGM attenuated line profiles with models lacking the IGM component in Sect. 4.10.
@@ 4.1 Individual reconstructed ${\rm Ly}\alpha$ line profiles
In Fig. 2, we show nine examples of the performance of zELDA in observed ${\rm Ly}\alpha$ line profiles. The observed ${\rm Ly}\alpha$ line profile is shown in grey, while the predictions of IGM+z, IGM-z, and NoIGM are shown in blue, green, and yellow. The top row shows the HST ${\rm Ly}\alpha$ line profiles (sources A, B and C) while the middle and bottom rows show MUSE LAEs (sources from D to I ).
@@
In general, IGM+z, IGM-z, and NoIGM reproduce the red peak of the observed ${\rm Ly}\alpha$ line profiles well. IGM+z and IGM-z predict a blue peak while differing from NoIGM. The NoIGM predicted ${\rm Ly}\alpha$ line profiles closely follow the observed line profile. This is the case for the HST and MUSE spectrum.  [^4] Note that, in mock ${\rm Ly}\alpha$ line profiles, a high $z$ observed line profile without a blue peak is not always reconstructed by IGM+z and IGM-z with a blue peak.
Figure 5: Rest frame difference between zELDA’s ${\rm Ly}\alpha$ wavelength and the estimated ${\rm Ly}\alpha$ wavelength using other spectral feature different from ${\rm Ly}\alpha$ in HST sources. zELDA’s predictions are shown blue, green and yellow for IGM+z, IGM-z and NoIGM, respectively. The mean and standard deviation of the distributions are shown for each model.
@@
At low redshift (top row of Fig. 2), the three ANN models (IGM+z, IGM-z, and NoIGM) predict nearly identical intrinsic line profiles for sources A and B, and the predicted profiles align closely with the observed spectra, particularly in both blue and red peaks. This consistency reflects the minimal IGM attenuation at these redshifts. In source C ($z=0.36$), all models predict similar red peak amplitudes, but the blue peak exhibits slight differences: IGM-z predicts the strongest blue peak, followed by IGM+z and then NoIGM. These differences correlate with variations in the predicted parameters, such as $V_{\rm exp}$, which decreases from 96.4 km s-1 (NoIGM) to 71.6 km s-1 (IGM+z) and 51.6 km s-1 (IGM-z). This trend indicates that the inclusion of IGM attenuation affects the reconstructed expansion velocity and suggests that IGM-z provides a more accurate fit by incorporating additional IGM constraints.
@@
At higher redshift (middle and bottom rows), the influence of IGM absorption becomes more pronounced. For sources D, E, G, H, and I, IGM+z and IGM-z consistently predict a reconstructed blue peak, albeit attenuated or partially absorbed, while NoIGM either predicts no blue peak or one with significantly lower amplitude. For instance, in source I, the suppressed blue peak predicted by IGM+z and IGM-z ($f_{\rm esc}^{4\AA}$ $\sim 0.56$) matches the observations, whereas NoIGM fails to replicate this feature, highlighting the necessity of IGM modeling in accurately capturing line profiles at high redshift. Furthermore, the small difference in $f_{\rm esc}^{4\AA}$ values (within 5%) between IGM+z and IGM-z underscores the robustness of their predictions.
@@
The observed and reconstructed ${\rm Ly}\alpha$ line profiles of all the sources are shown in Appendix C. Fig. 18 for sources with $0.200.0<z<$ 0.20, Fig. 19 for sources with $0.350.20<z<$ 0.35, Fig. 20 for sources with $3.30.35<z<$ 3.3 , Fig. 21 for sources with $3.73.3<z<$ 3.7 , Fig. 22 for sources with $4.23.7<z<$ 4.2 ,Fig. 23 for sources with $4.84.2<z<$ 4.8 , Fig. 24 for sources with $6.04.8<z<$ 6.0. In general, the red peak of the ${\rm Ly}\alpha$ predicted by the ANN models follows the observations closely. The blue peak predictions follow the observations at low redshift, while at high redshift, they diverge for IGM+z and IGM-z.
4.2 Comparison of Outflow Properties across Models
In Fig. 4, we show a comparison between the output variables of IGM+z (X-axis) and those predicted by NoIGM (Y-axis). In general we find that there is a Spearman correlation coefficient higher than $\sim 0.7$ for the whole sample. The Spearman correlation coefficient is greater in the low redshift sample than in the high redshift sample. For example, for $V_{\rm exp}$, $\rho_{z<3}=0.72$ while $\rho_{z>3}=0.69$ and for $N_{\rm H}$ $\rho_{z<3}=0.83$ while $\rho_{z>3}=0.64$. The IGM+z model predicts higher values for $V_{\rm exp}$ at high redshift. Meanwhile, IGM+z predicts lower values of $W_{\rm in}$ than NoIGM at high redshift.
Figure 6: Left: Distribution of the Kolmogórov-Smirnov (KS) estimator in the red tail of HST ${\rm Ly}\alpha$ line profiles for IGM+z, IGM-z, and NoIGM in thick blue, medium green, and narrow yellow, respectively. The width of the distribution is shown for each distribution, computed as the difference between the percentiles 84 and 16, and the median.
Figure 7: Comparison of the ANN output and LASD variables in the observed ${\rm Ly}\alpha$ line profiles predicted by the same models. $EW_{\rm obs}$ and $L_{\rm Ly\alpha}$ were obtained from LASD. $V_{\rm exp}$ is given in  km s-1, $N_{\rm H}$ in $cm^{-2}$, $EW_{\rm in}$, $EW_{\rm ob}$ and $W_{\rm in}$ in Å and $L_{\rm Ly\alpha}$ in $erg/s$. The blue dots show the values predicted by IGM+z while the green dots show that of IGM-z. The grey dots show the properties that include measurements by LASD. The histograms on the diagonal show the distribution of the parameters in the corresponding color for each model.
@@ 4.2 Comparison of Outflow Properties across Models
4.3 Accuracy of redshift estimates with zELDA
In this section, we study the redshift accuracy of zELDA. For this goal, we make use of the HST sources with a systemic redshift measured with a different feature than ${\rm Ly}\alpha$. We focus on HST sources as MUSE sources in LASD lack a systemic redshift below 0.5. This makes a total of 111 sources. To quantify the precision of the redshift, we analyze the distribution of the difference in rest frame wavelength between the predicted redshift by zELDA $z^{\rm zELDA}$ and the systemic redshift $z^{\rm Sys}$, computed as in Gurung-López et al. (2022), i.e.,
shown in Fig. 5. We find that zELDA provides a sub-angstrom accuracy using only the ${\rm Ly}\alpha$ line profile.
Our analysis reveals that both IGM+z and IGM-z outperform NoIGM in estimating the redshift of the sources, with IGM-z showing slightly lower dispersion compared to IGM+z. The comparison highlights the robustness of the IGM-inclusive models in reducing uncertainty and providing consistent predictions.
However, a small systematic bias is present across all models, with zELDA generally overestimating the ${\rm Ly}\alpha$ rest frame by approximately 0.25Å (equivalent to  61 km s-1). This overestimation aligns with findings from previous studies, such as Orlitová et al. (2018) and Li et al. (2021), and points to inherent limitations of the ‘thin shell model.’
4.4 Line profile fitting accuracy
To compare the models fairly, the accuracy of the recovered shape of the intrinsic ${\rm Ly}\alpha$ line profile in the observed spectrum, we focus on their red tails. In principle, galaxies at low redshift should be little absorbed by the IGM. In particular, most of the absorption should happen bluewards ${\rm Ly}\alpha$, leaving mostly untouched the red side of ${\rm Ly}\alpha$. Thus, the peak of the red peak and its tail extending redwards should be close or identical to that of the ${\rm Ly}\alpha$ line profile before traveling through the IGM. Therefore, comparing the red tail of the observed ${\rm Ly}\alpha$ line profiles to the predicted intrinsic spectrum in the same wavelength range should be a fair comparison. We define the red tail of our ${\rm Ly}\alpha$ line profiles as the wavelength range between the maximum of the red peak and ${\rm Ly}\alpha$ +5Å.
In Fig. 6, we show the distribution of the KS estimator for the IGM+z (blue), IGM-z (green), and NoIGM (yellow) in the red tails of the HST sources. We find that IGM-z and IGM+z outperform NoIGM. In particular, the median of the KS distribution for the models are 0.07 for IGM+z, 0.07 for IGM-z and 0.09 for NoIGM. Meanwhile, NoIGM exhibit a little bit more dispersion than IGM+z and IGM-z. We conclude that the IGM+z and IGM-z models predict the red tail of the observed ${\rm Ly}\alpha$ line profile with the same precision, being IGM-z only slightly better than IGM+z.
Figure 10: Mean $f_{\rm esc}^{4\AA}$ as a function of redshift (squares) for IGM+z (blue) and IGM-z (green). The solid lines indicate the fits, including all the sources. Meanwhile, the colored dashed line indicates the fit using only MUSE data. The grey dashed line shows the Faucher-Giguère et al. (2008) mean IGM transmission.
@@ 4.4 Line profile fitting accuracy
0.0-5.0 HST+MUSE IGM+$z$ 5.35$\pm$0.0143 0.8$\pm$0.0084 0.93$\pm$0.0002
0.0-5.0 HST+MUSE IGM-$z$ 5.5$\pm$0.0366 0.86$\pm$0.0196 0.89$\pm$0.0002
2.9-5.0 MUSE IGM+$z$ 5.18$\pm$0.187 0.66$\pm$0.084 1.0$\pm$0.033
2.9-5.0 MUSE IGM-$z$ 5.25$\pm$0.326 0.66$\pm$0.133 1.0$\pm$0.057
4.5 Correlations between model and observed properties
In this section, we study the existence of a correlation between the thin shell outflow properties predicted by zELDA, redshift and the ${\rm Ly}\alpha$ luminosity $L_{\rm Ly\alpha}$ and rest frame observed ${\rm Ly}\alpha$ equivalent width $EW_{\rm ob}$. The latter two quantities are obtained directly from LASD.
In Fig. 7, we show the predicted properties against all the other properties. $V_{\rm exp}$, $N_{\rm H}$, $\tau_{a}$, $EW_{\rm in}$, $W_{\rm in}$, $z$, $f_{\rm esc}^{4\AA}$, $EW_{\rm ob}$ and $L_{\rm Ly\alpha}$ from left to right and from top to bottom. The top right triangle ( green and grey) shows the properties predicted by IGM-z against themselves. The bottom left triangle (blue and grey) shows those of IGM+z. The diagonal shows the distribution of the parameters.
Meanwhile, the Spearman correlation coefficient for the entire sample is shown in Fig. 8 with the $1\sigma$ uncertainty. The top left corner shows the Spearman correlation coefficient for IGM-z, $\rho_{-z}$(green). Meanwhile, the bottom left corner shows the Spearman correlation coefficient for IGM+z, $\rho_{+z}$(blue). To estimate the uncertainty proceed as in Gurung-López et al. (2022), we draw 1000 times a new set of outflow parameters from the posterior of the ANN models (see the methodology in ZP24). We compute the Spearman correlation coefficient for each set of parameters. The $1\sigma$ uncertainty is calculated by the percentiles 16 and 84 from the distribution of Spearman coefficients. We consider that two properties correlate or anti-correlate when the absolute value of the Spearman correlation coefficient is greater than 0.45. Furthermore, the Spearman correlation coefficients for the MUSE and HST separately are presented in Appendix B.
We find that, in general, the Spearman coefficient provided by the IGM+z and IGM-z models are similar. Typically, both models exhibit or lack the same correlations. Also, Fig. 17 shows the Spearman correlation coefficient for the low redshift (HST, left) and high redshift samples (MUSE right). Considering the complete ${\rm Ly}\alpha$ line profile sample, we find that only a handful of properties correlate. Overall, some of these correlations are also present in MUSE and HST samples by themselves, but others are missing.
We find the following relations:
@@ 4.6 IGM escape fraction evolution with redshift
Figure 11: Redshift evolution of different ${\rm Ly}\alpha$ escape fractions. The filled colored squares show the $\langle f_{\rm esc}^{4\AA}\rangle$ found by the IGM+z and IGM-z models up to $z=5.0$, while those empty show the results at $z>5$. The colored solid lines show the best fit to $\langle f_{\rm esc}^{4\AA}\rangle$ as presented previously. The dashed red line shows the global ${\rm Ly}\alpha$ escape fraction estimation by Hayes et al. (2011). The grey symbols show different studies from the literature measuring the global ${\rm Ly}\alpha$ escape fraction and their uncertainty. The colored dashed line shows the ISM escape fraction estimated from the combination of Hayes et al. (2011) and our results.
In this section, we explore in detail the $f_{\rm esc}^{4\AA}$ redshift dependence predicted by the models including an IGM component (igmz and IGM-z).
As shown in ZP24 (section 4.3 and appendix C), the accuracy of the IGM escape fraction $f_{\rm esc}^{4\AA}$ depends on the spectral quality of the ${\rm Ly}\alpha$ line profiles. In this section, we restrain our analysis to ${\rm Ly}\alpha$ line profiles with $S/N_{p}$8.0 and $W_{\rm g}$3.0Å. This results in a sample composed of 93 HST and 113 MUSE sources, totaling 206 ${\rm Ly}\alpha$ line profiles.
In Fig. 9, we show the $f_{\rm esc}^{4\AA}$ values for the 206 sources given by IGM+z (left) and IGM-z (middle), as well as their distributions (right). We find that $f_{\rm esc}^{4\AA}$ decreases with increasing redshift in both ANN models. In particular, $f_{\rm esc}^{4\AA}$ in HST sources ($z<0.55$) is concentrated between 0.8 and 1.0 for both IGM+z and IGM-z. Considering MUSE sources at $z>3.0$, $f_{\rm esc}^{4\AA}$ is concentrated between 0.85 and 0.5 for both IGM+z and IGM-z with dispersion $\sim 0.2$.
This intuitively makes sense, of course, as the higher neutral hydrogen content and increased IGM density at greater redshifts lead to stronger absorption of ${\rm Ly}\alpha$ photons, thereby reducing the IGM escape fraction.
In Fig. 10, we show the mean ${\rm Ly}\alpha$ IGM escape fraction, $\langle f_{\rm esc}^{4\AA}\rangle$ for the models IGM-z (blue squares) and IGM+z (green squares). The uncertainties were computed as the uncertainty of the mean. The bins have been shifted $\pm$0.01 in redshift for displaying proposes. We find that the $\langle f_{\rm esc}^{4\AA}\rangle$ predicted by IGM+z and IGM-z diverge at $z<0.5$ in HST data, while they match at $z>3.0$ in MUSE data. In HST data, the IGM+z and the IGM-z models predict no evolution in $\langle f_{\rm esc}^{4\AA}\rangle$ with redshift. IGM+z predicts a constant value of $\sim$0.91, while IGM-z predicts $\sim$0.89. Meanwhile, at $z>3.0$, both models predict that $\langle f_{\rm esc}^{4\AA}\rangle$ changes from 0.9 at $z=$2.9 to 0.6 at $z=5.0$.
In Fig. 10 we also show the best fitting of the predicted $\langle f_{\rm esc}^{4\AA}\rangle$ to a modulated Fermi-Dirac distribution (zELDA is able to recover the $f_{\rm esc}^{4\AA}$-$z$ evolution as discussed in ZP24), i.e.,
@@
The best fitting parameters can be found in Tab.1. We find that fitting the whole redshift range (HST+MUSE) for IGM+z gives a best fitting $c=0.93$, while for IGM-z $c=0.89$. . However, in the $5.02.9<z<$ 5.0 range, the parametric curves are closer than than 0.05. Interestingly, when fitting only MUSE data ($5.02.9<z<$ 5.0), the best fitting curve for both ANN models are very similar. In fact, for both models $c=1.0$. Also, $a$ and $b$ are $1\sigma$ compatible. This causes a difference at $z<2.5$ in the best-fitting curves when considering the complete and only MUSE samples
For comparison, we also show $\langle T_{\rm IGM}\rangle$, the mean IGM transmission (Faucher-Giguère et al., 2008) in grey. The relationship between $\langle T_{\rm IGM}\rangle$ and $f_{\rm esc}^{4\AA}$ is a priory unknown: $\langle T_{\rm IGM}\rangle$ reflects the mean IGM transmission. In contrast, $f_{\rm esc}^{4\AA}$ measures the escape of ${\rm Ly}\alpha$ from LAEs on intermediate IGM scales and the ${\rm Ly}\alpha$ line profiles emerging from the ISM and CGM. For instance, if red peak dominated ${\rm Ly}\alpha$ line profiles enter into the IGM, the ${\rm Ly}\alpha$ IGM escape fraction $f_{\rm esc}^{4\AA}$ will be high compared to blue peak dominated ${\rm Ly}\alpha$ line profiles. $\langle f_{\rm esc}^{4\AA}\rangle$ predicted by IGM-z and IGM+z fall 0.1 below $\langle T_{\rm IGM}\rangle$. The difference between $f_{\rm esc}^{4\AA}$ and $\langle T_{\rm IGM}\rangle$ can be due to the stronger absorption around the ${\rm Ly}\alpha$ line center (Laursen et al., 2011b; Gurung-López et al., 2022). Even if the IGM transmission is $\sim 1$ a few Angstroms from ${\rm Ly}\alpha$, the stronger absorption at $-2\AA{}<\Delta\lambda<0.5\AA{}$ can affect the blue peak of $z=0$ lines. Yet, at high redshift ($z>1.5$), both, the parametric best fitting curves for both ANN models and the predicted $\langle f_{\rm esc}^{4\AA}\rangle$ are above $\langle T_{\rm IGM}\rangle$. While implying a non-trivial $\langle f_{\rm esc}^{4\AA}\rangle$-$\langle T_{\rm IGM}\rangle$ relationship evolution, this trend could also indicate a selection effect. In this work, we measure $f_{\rm esc}^{4\AA}$ in MUSE galaxies with observed strong ${\rm Ly}\alpha$ emission rather than the whole underlying galaxy population. Hence, it is likely that measuring $f_{\rm esc}^{4\AA}$ in ${\rm Ly}\alpha$ flux-limited samples might result in $\langle f_{\rm esc}^{4\AA}\rangle$ values above those of the whole star-forming galaxy population.
4.7 Comparison between the global and IGM escape fractions
The global ${\rm Ly}\alpha$ escape fraction, $f_{\rm esc}^{\rm Lya}$, is defined as the ratio between the total amount of ${\rm Ly}\alpha$ photons produced inside a galaxy and those reaching the observer. Thus, $f_{\rm esc}^{\rm Lya}$ encapsulates information about the escape of ${\rm Ly}\alpha$ from the ISM, CGM, and IGM. The redshift evolution of $f_{\rm esc}^{\rm Lya}$ has been explored in the literature (Blanc et al., 2011; Ouchi et al., 2008; Cassata et al., 2011; Sobral et al., 2018; de La Vieuville et al., 2019; Thai et al., 2023; Goovaerts et al., 2024). In particular, Hayes et al. (2011) calculated the $f_{\rm esc}^{\rm Lya}$ redshift evolution comparing the ${\rm H}\alpha$ flux, mostly unaffected by radiative transfer processes, to the measured ${\rm Ly}\alpha$ flux in the same sources.
In Fig. 11, we show the $f_{\rm esc}^{\rm Lya}$ redshift evolution obtained by Hayes et al. (2011) (dashed red line) along with the individual data points from the other studies in grey as described in the legend. Other studies, such as Goovaerts et al. (2024) (based on Thai et al. (2023)), also obtained relations between the redshift and $f_{\rm esc}^{\rm Lya}$ , finding compatible results with Hayes et al. (2011). Overall, it is found in the literature that $f_{\rm esc}^{\rm Lya}$ increases with redshift, being earlier galaxies less attenuated in ${\rm Ly}\alpha$. Normally, this is attributed to the dust in the ISM of galaxies, which is less abundant at high redshift Hayes et al. (2011); Gurung-López et al. (2019b); Goovaerts et al. (2024). In particular, the different experiments find that $f_{\rm esc}^{\rm Lya}<5\%$ at $z<3.5$. Also, at $65<z<$ 6, Ouchi et al. (2008); Cassata et al. (2011); de La Vieuville et al. (2019) found a $f_{\rm esc}^{\rm Lya}$ bellow 0.2 with relatively low uncertainty of $\sim 0.1$. Meanwhile, with lager uncertainties, Sobral et al. (2018) found a $f_{\rm esc}^{\rm Lya}\sim 0.3$ at $64<z<$ 6 and Goovaerts et al. (2024) found $f_{\rm esc}^{\rm Lya}\sim 0.8$ at $z=6$. Given the uncertainties, all these studies are compatible with each other.

### v2

@@ 4 Results on observed Lyman-$\alpha$ line profiles
In this section, we apply zELDA’s ANN models to observed ${\rm Ly}\alpha$ line profiles. In Sect. 4.1 we present a few examples of the predicted ${\rm Ly}\alpha$ line profiles. We study the correlation between $f_{\rm esc}^{4\AA }$ and outflow properties in Sect. 4.2. We present our observed $\langle f_{\rm esc}^{4\AA }\rangle$ reconstruction in Sect. 4.3.
@@ 4.1 Individual reconstructed ${\rm Ly}\alpha$ line profiles
In Fig. 2, we present nine examples illustrating the performance of zELDA on observed ${\rm Ly}\alpha$ line profiles. The observed spectra are shown in grey, while the predictions of IGM+z, IGM-z, and NoIGM are displayed in blue, green, and yellow, respectively. The top row corresponds to HST sources (A, B, and C), whereas the middle and bottom rows show MUSE LAEs (DI).
@@
Overall, the three ANN models reproduce the red peak of the observed ${\rm Ly}\alpha$ line profiles well. Both IGM+z and IGM-z also predict the presence of a blue peak, while NoIGM tends to follow the observed spectra more closely without introducing additional features. This behavior is consistent across both HST and MUSE data. [^6] Note that in mock ${\rm Ly}\alpha$ line profiles, high-$z$ spectra lacking an observed blue peak are not always reconstructed by IGM+z and IGM-z with a blue peak.
@@
At low redshift (top row of Fig. 2), the three models predict nearly identical intrinsic profiles for sources A and B, with close agreement between the predicted and observed spectra in both the red and blue peaks. This consistency reflects the negligible IGM attenuation at these redshifts. For source C ($z=0.36$), all models produce similar red peak shapes and amplitudes, but differ slightly in the blue peak: IGM-z predicts the strongest blue component, followed by IGM+z and then NoIGM. These differences correlate with the reconstructed expansion velocity, which decreases from 96.4 km s-1 (NoIGM) to 71.6 km s-1 (IGM+z) and 51.6 km s-1 (IGM-z). This trend suggests that including IGM attenuation impacts the inferred kinematic parameters, with IGM-z providing a more accurate fit by incorporating IGM constraints.
@@
At higher redshift (middle and bottom rows), IGM absorption becomes more significant. For sources D, E, G, H, and I, both IGM+z and IGM-z predict a reconstructed blue peak—attenuated or partially absorbed—whereas NoIGM either predicts no blue peak or one with much lower amplitude. For instance, in source I, the suppressed blue peak predicted by IGM+z and IGM-z ($f_{\rm esc}^{4\AA }$ $\sim 0.56$) matches the observations, while NoIGM fails to reproduce this feature. This highlights the importance of IGM modeling for accurately reconstructing line profiles at high redshift. Moreover, the small difference in $f_{\rm esc}^{4\AA }$ values (within 5%) between IGM+z and IGM-z further underscores the robustness of their predictions.
@@
The observed and reconstructed ${\rm Ly}\alpha$ line profiles of all the sources are shown in Appendix F. Fig. 19 for sources with $0.0<z<0.20$, Fig. 20 for sources with $0.20<z<0.35$, Fig. 21 for sources with $0.35<z<3.3$ , Fig. 22 for sources with $3.3<z<3.7$ , Fig. 23 for sources with $3.7<z<4.2$ ,Fig. 24 for sources with $4.2<z<4.8$ , Fig. 25 for sources with $4.8<z<6.0$. In general, the red peak of the ${\rm Ly}\alpha$ predicted by the ANN models follows the observations closely. The blue peak predictions follow the observations at low redshift, while at high redshift, they diverge for IGM+z and IGM-z.
4.2 Correlations between model and observed properties
@@
We quantify the strength of the possible correlations by the Spearman correlation coefficients for the full sample, which are shown in Fig. 3, along with their $1\sigma$ uncertainties. Meanwhile, the full 1-dimensional and 2-dimensional distributions can be found in Appendix A. The coefficients from IGM-z are shown in the top-right corner (green, $\rho$), and those from IGM+z in the bottom-left corner (blue, $\rho_{z}$). Uncertainties are estimated following Gurung-López et al. (2022): we draw 1000 realizations of the outflow parameters from the ANN posteriors (see methodology in Z25), compute the Spearman correlation coefficient for each realization, and define the $1\sigma$ interval from the 16th and 84th percentiles of the resulting distribution. We consider two properties to be correlated (or anti-correlated) if the absolute value of the Spearman coefficient exceeds 0.45. Correlation matrices for the HST and MUSE subsamples separately are provided in Appendix D.
Overall, we find that the correlation coefficients inferred from IGM+z and IGM-z are very similar: both models typically identify the same correlations. In Appendix D, we present the coefficients for the low-redshift (HST, left) and high-redshift (MUSE, right) samples (see Fig. 17). Across the full dataset, only a handful of parameters show significant correlations, some of which are present in both subsamples (low and high redshift), while others appear only in one. These relations are the following:
[…]
Figure 5: Mean $f_{\rm esc}^{4\AA }$ as a function of redshift (squares) for IGM+z (blue) and IGM-z (green). The solid lines indicate the fits, including all the sources below $z=5$. Meanwhile, the colored dashed line indicates the fit using only MUSE data below $z=5$.
@@ 4.3 IGM escape fraction evolution with redshift
In this section, we examine in detail the redshift dependence of the IGM escape fraction, $f_{\rm esc}^{4\AA }$, as predicted by the models that include an IGM component (IGM+z and IGM-z).
As shown in Z25 (Sect. 4.3 and Appendix C), the accuracy of $f_{\rm esc}^{4\AA }$ depends strongly on the spectral quality of the ${\rm Ly}\alpha$ line profiles. We therefore restrict our analysis to spectra with $S/N_{p}$ >8.0>8.0 and $W_{\rm g}$ <3.0<3.0Å. This selection yields 93 HST and 113 MUSE sources, for a total of 206 ${\rm Ly}\alpha$ line profiles. These cuts ensure that the uncertainty in $f_{\rm esc}^{4\AA }$ remains below 0.05 for HST sources and below 0.09 for MUSE sources.
[…]
In Fig. 5, we present the mean ${\rm Ly}\alpha$ emission line IGM escape fraction, $\langle f_{\rm esc}^{4\AA }\rangle$, for IGM-z (blue squares) and IGM+z (green squares). Uncertainties correspond to the error of the mean, and the bins have been shifted by $\pm 0.01$ in redshift for clarity. At $z<0.5$ (HST data), IGM+z and IGM-z diverge slightly, with IGM+z predicting a constant value of $\sim 0.91$ and IGM-z $\sim 0.89$. At $z>3.0$ (MUSE data), however, both models agree closely, predicting that $\langle f_{\rm esc}^{4\AA }\rangle$ declines from $\sim 0.9$ at $z=2.9$ to $\sim 0.6$ at $z=5.0$.
In Fig. 5 we also show the best fitting of the predicted $\langle f_{\rm esc}^{4\AA }\rangle$ to a modulated Fermi-Dirac distribution (zELDA is able to recover the $f_{\rm esc}^{4\AA }$-$z$ evolution as discussed in Z25), i.e.,
@@
The best-fitting parameters are listed in Tab. 1. For the full redshift range (HST+MUSE), the fits yield $c=0.93$ for IGM+z and $c=0.89$ for IGM-z. In the interval $2.9<z<5.0$, the parametric curves differ by less than 0.05. When restricted to MUSE data alone, both ANN models produce nearly identical fits, with $c=1.0$ and $a$ and $b$ consistent within $1\sigma$. This explains the differences at $z<2.0$ between the best-fitting curves derived from the full sample and from MUSE-only data. Extrapolating the fits, both models predict that the ${\rm Ly}\alpha$ IGM escape fraction approaches zero at $z\sim 11$.
[…]
5 Validation and discussion
In this section we aim to validate our results on the reconstructed ${\rm Ly}\alpha$ line profile shapes and $\langle f_{\rm esc}^{4\AA }\rangle$ redshift evolution. This task is made complex by an inherent aspect of Lya line profile reconstruction: we can only observe one sight line per source. In Z25 we calibrated and tested zELDA using multiple IGM transmission sight lines for the same ${\rm Ly}\alpha$ line profile, which is possible only for simulated spectra. We refer the reader to Z25 for an in-depth analysis of the performance of zELDA on mock ${\rm Ly}\alpha$ spectra with HST and MUSE quality. In order to show the robustness of the ${\rm Ly}\alpha$ line profile reconstruction and $\langle f_{\rm esc}^{4\AA }\rangle$ estimation we performed the following tests:
  • We show that the inferred redshift is highly accurate ($\sim$0.3Å) , with IGM-z being the least biased (see Sect. 5.1).
  • We show that the red tail of the ${\rm Ly}\alpha$ line profile, supposedly unaffected by the IGM at low redshift, is well recovered, specially when implementing the IGM absorption (see Sect. 5.2).
  • We computed the ${\rm Ly}\alpha$ stacked line profiles from the reconstructed ${\rm Ly}\alpha$ line profiles (see Sect. 5.3). We find that the stacked ${\rm Ly}\alpha$ line profiles do not evolve strongly with redshift. This indicates that the evolution in the observed ${\rm Ly}\alpha$ spectrum arises from the IGM absorption, in agreement with previous findings in the literature (Hayes et al., 2021a, 2023).
  • We demonstrate that our $\langle f_{\rm esc}^{4\AA }\rangle$ measurement in individual ${\rm Ly}\alpha$ line profiles is compatible with the IGM mean optical depth found in the literature, especially at $z>2$ (see Sect. 5.4) .
  • We find that our measurement of the IGM ${\rm Ly}\alpha$ escape fraction is consistent with the global ${\rm Ly}\alpha$ escape fraction reported in the literature in the whole redshift range (see Sect. 5.5).
Finally, we discuss the fact that the IGM+z and IGM-z models predict that the ${\rm Ly}\alpha$ line profiles are IGM attenuated at $z\sim 0$ in Sect. 5.6.
5.1 Accuracy of redshift estimates with zELDA
In this section, we evaluate the redshift accuracy of zELDA. For this purpose, we use HST sources with systemic redshifts measured from features other than ${\rm Ly}\alpha$, yielding a total of 111 sources. We focus on HST data because MUSE sources in LASD lack systemic redshift measurements below $z=0.5$. To quantify the precision, we analyze the distribution of rest-frame wavelength differences between the redshift predicted by zELDA ($z^{\rm zELDA}$) and the systemic redshift ($z^{\rm Sys}$), computed as in Gurung-López et al. (2022):
[…]
Our analysis shows that the three models IGM+z and IGM-z and NoIGM perform similarly. All of them recover accurately the ${\rm Ly}\alpha$ wavelength location in the line profiles with an uncertainty of $\sim 0.3$Å. At low redshift, where the IGM has only a minor impact on the ${\rm Ly}\alpha$ line profile, the performance gain of IGM+z and IGM-z over NoIGM is modest. However, at higher redshifts the effect of the IGM becomes more significant, and we expect the IGM-inclusive models to substantially outperform NoIGM.
A small systematic bias is present across all models, with zELDA generally overestimating the ${\rm Ly}\alpha$ rest frame by approximately 0.25Å (equivalent to $\sim$61 km s-1). This overestimation aligns with findings from previous studies, such as Orlitová et al. (2018) and Li et al. (2021), and points to inherent limitations of the ‘thin shell model.’ It is noticeable that IGM+z and NoIGM are biased evenly while IGM-z shows a better performance.
5.2 Line profile fitting accuracy
Figure 6: Rest frame difference between zELDA’s ${\rm Ly}\alpha$ wavelength and the estimated ${\rm Ly}\alpha$ wavelength using spectral features different from ${\rm Ly}\alpha$ in HST sources. zELDA’s predictions are shown blue, green, and yellow for IGM+z, IGM-z, and NoIGM, respectively. The mean and standard deviation of the distributions are shown for each model.
Figure 7: Left: Distribution of the Kolmogórov-Smirnov (KS) estimator in the red tail of HST ${\rm Ly}\alpha$ line profiles for IGM+z, IGM-z, and NoIGM in thick blue, medium green, and narrow yellow, respectively. The width of the distribution is shown for each distribution, computed as the difference between the 84th and 16th percentiles and the median.